OpenStudy (anonymous):

In a circle of diameter 40cm the lenght of a chord is 20cm. Find the length of the major arc corresponding to the chord.

OpenStudy (anonymous):

[(100pi)/3] cm?

OpenStudy (anonymous):

@eliassaab A minor mistake \[2\Pi-(\Pi/3)=5\Pi/3\]

OpenStudy (anonymous):

Old age. I do not know how to subtractt anymmore.

OpenStudy (anonymous):

Ha Ha... happens

OpenStudy (anonymous):

Thanks for @scsam here is a correction; The circle formed by the center in the two endpoints of the chord is equilateral. So the small arc is \[ \frac \pi 3 \] and the big chord is \[ 2\pi -\frac \pi 3 = \frac{5\pi} 3 \] So the length of it is \[20 \frac{5\pi} 3 =\frac {100 \pi}{3} \]